OpenStudy (anonymous):

TRIGONOMETRY... HELP FOR MEDAL? If sin Θ = 2/3 and tan Θ < 0, what is the value of cos Θ?

OpenStudy (anonymous):

Since sine =y/r, y=2 and r=3 Then x= sqrt[3^2-2^2] sin(t) is positive and tan(t) is negative, t is the QII where x is negative a y is positive. cos(t)= x/r= -sqrt(5)/3

OpenStudy (anonymous):

Thank you so much! Can you help me with 2 more?

OpenStudy (anonymous):

Send me the answers

OpenStudy (anonymous):

Nvm I got a 100! Thanks so much. :)